# A novel sensitivity analysis method for agent-based models stratifies in-silico tumor spheroid simulations

### v2

@@ 4 Discussion and conclusions
[…]
@@
Identify common yes no yes
@@
Computationally efficient yes no further investigation required
@@
Another strength of SSRCA is its flexibility with the choice of methodologies used within the pipeline. We found similar results in this work whether we used the Cell counts DV or Final Cell Density DV. This robustness to DV choice is crucial, as ABMs are broadly used for biological problems, and the appropriate DV choices depends on the application. For example, many studies rely on order parameters to characterize ABM behavior; alternatively, topological data analysis (TDA) is a recent field of research that has found success in summarizing many types of ABMs with applications in complex systems [36, 53, 54]. SSRCA can be used with any of these DVs, as it only requires they are uniform length. Additionally, we used principal components analysis and $k$-means clustering for dimensionality reduction and clustering, respectively, in this work, but any suitable algorithm can be used for these processes.
@@
We applied the SSRCA methodology to a 2-dimsensional ABM of tumor spheroid growth and identified $c_{a}$, $c_{d}$, $\eta_{1}$, $\eta_{3}$ as the most sensitive parameters. This result suggests that the most impactful mechanisms for tumor spheroid phenotype are cell cycle entry and cell death, which may inform the future use of tumor spheroid models to interpret experimental data. For example, 3-dimsensional Klowss Model was calibrated to experimental data from human primary melanoma cells [21]; modeling similar data from other cell lines will likely require different parameter values. Our work proposes that thoroughly exploring all unknown parameters is not necessary; instead, one may more efficiently vary the 4 identified sensitive parameters to match the model to new data. Our proposed methodology may be of use for other modeling frameworks. For example, the authors of [55] were could not perform parameter estimation due to their differential equation model containing 19 unknown parameters. The SSRCA methodology could be applied to this model to learn an informative subset of its parameters for parameter estimation.
@@
There are several avenues for future work with the SSRCA algorithm. While we can use the SSRCA Methodology to identify sensitive parameters, it is currently not able to rank parameters based on their sensitivity. We plan to further analyze the model and/or develop metrics that can be used in combination with SSRCA to enable such parameter ranking. We also demonstrated the utility of the SSRCA Method when applied to informative DVs, however, it has been reported previously that some DVs are more informative than others depending on the context [36, 34, 56, 57]. It would be interesting to explore if the SSRCA algorithm can be used to compare the information content between various DVS, or even to propose an optimal DV choice. We applied the SSRCA Methodology to a simple ABM of tumor spheroid growth in this work, but we plan to use it to interpret more complex models in the future, including models for other biological contexts, including cell biology, epidemiology, and ecology.
Data and code availability: All code and simulated data for this work is publicly available at
https://github.com/e-rohr/FUCCI_ABM2D.
@@
Appendix A Model overview
We simulate an agent-based model of tumor spheroid growth developed in [21]. This model is hybrid, meaning we track both discrete agent locations and a continuous nutrient concentration (e.g., oxygen) within the spheroid. Tumor cells are represented as discrete agents centered at $\textbf{x}_{n}(t)=(x_{n}(t),y_{n}(t))$ for $n=1,\ldots,N(t)$, where $N(t)$ is the total number of agents at time $t$. The spatio-temporal distribution of the nutrient concentration is non-dimensionalized and represented by $c(\textbf{x},t)\in[0,1]$. In addition to agent locations, we also model how cells progress through the cell cycle. The agents are colored according to FUCCI: agents in the G1 phase are colored red, agents in the early S (eS) phase are colored yellow, and agents in the G2/S/M phase are colored green. Following [21], we assume the non-dimensional, local nutrient concentration, $c(\textbf{x},t)$, regulates cell motility, proliferation, and death [8, 58, 10]. The model is simulated in a cubic domain, $\Omega$, with side length $1000~{}\mu\mathrm{m}$, which we chose to mitigate computational costs while ensuring that agents do not reach the domain boundary during simulation.
We now describe how tumor cells in the model spheroid undergo mitosis, cell motility, and cell death, and then detail how the nutrient spreads and is consumed. All model parameters are described in Table 1. The tumor cell processes are modelled as discrete events using the Gillespie Algorithm [59]. We initialize the ABM in a similar manner as [21], where initial cell colors are randomized to produce proportions of red, yellow, and green cells that match experimental observations.
A.1 Agent Dynamics
@@
$m(c)$ $=(m_{max}-m_{min})\frac{c^{\eta_{2}}}{c_{m}^{\eta_{2}}+c^{\eta_{2}}}+m_{min},$ (6)
$d(c)$ $=(d_{max}-d_{min})\left(1-\frac{c^{\eta_{3}}}{c_{d}^{\eta_{3}}+c^{\eta_{3}}}\right)+d_{min},$ (7)
where all parameters are described in Table 1. When migrating, an agent is displaced $\mu$ units in the randomly chosen direction. When an agent dies, it is removed from the simulation by setting its status to dead, and its final location is recorded.
A.2 Nutrient dynamics
@@
$\frac{\partial C}{\partial t}=D\nabla^{2}C-\kappa Cv,\text{ in }\Omega,$ (8)
@@
with diffusivity $D>0~{}[\mu\mathrm{m}^{2}~{}\mathrm{h}^{-1}]$, consumption rate $\kappa>0~{}[\mu m^{2}~{}(\mathrm{h}~{}\text{cells})^{-1}]$, and where $v(\textbf{x},t)\geq 0~{}[\text{cells }\mu m^{-2}]$ is the density of agents at position x and time $t.$ On the boundary, $\partial\Omega$, a maximum far-field concentration $C(\textbf{x},t)=c_{b}$ is imposed. We non-dimensionalize this independent variable by setting $c(\textbf{x},t)=C(\textbf{x},t)/c_{b}$, so that $c=1$ on $\partial\Omega$, and $c(\textbf{x},t)\in[0,1]$. We approximate the reaction-diffusion Equation (8) using the quasi-steady state equation
$0=\nabla^{2}c-\alpha cv,\text{ in }\Omega,$ (9)
@@
where $\alpha=\kappa/D>0\ [\mu\mathrm{m}~{}\mathrm{cell}^{-1}]$. During simulations, the spatio-temporal distribution of the chemical is modelled by solving Equation (9) $t^{*}$ units using a finite volume method on a uniform structured mesh with node spacing $h$. Following [21], we use $t^{*}=1$. The cell density $v(\textbf{x},t)$ is approximated by setting $v(\textbf{x}_{i,j},t)=N_{i,j}/h^{2}$, where $N_{i,j}$ is the number of agents within the control volume surrounding the node located at $(x_{i},y_{j})$ and $h^{2}$ is the volume of the control volume.
Figure 7: Determining the baseline $N(0)$ value. To determine the baseline value of $N(0)$, we visually compared simulations of the 2-dimensional Klowss model (top 5 rows) for various values of $N(0)$ to the baseline simulation of the 3-dimensional Klowss Model from [21] (bottom row). The latter depicts the $z=0$ cross section of the 3-dimensional simulation. The red text label indicates the parameter value we determined most closely matches the bottom row. The bottom row is reproduced and modified from [21], Copyright Klowss et al., and licensed under CC-BY 4.0 https://creativecommons.org/licenses/by/4.0/.
Figure 8: Determining the baseline $\alpha$ value. To determine the baseline value of $\alpha$, we visually compared simulations of the 2-dimensional Klowss model (top 5 rows) for various values of $\alpha$ to the baseline simulation of the 3-dimensional Klowss Model from [21] (bottom row). The latter depicts the $z=0$ cross section of the 3-dimensional simulation. The red text label indicates the parameter value we determined most closely matches the bottom row. The bottom row is reproduced and modified from [21], Copyright Klowss et al., and licensed under CC-BY 4.0 https://creativecommons.org/licenses/by/4.0/.
@@
Figure 9: Scree plots. The scree plots that result when performing dimensionality reduction using (a) Dataset 1 using the Cell counts DRDV, (b) Dataset 1 using the Final cell density DRDV, (c) Dataset 2 using the Cell counts DRDV, (d) Dataset 2 using the Final cell density DRDV
@@
Figure 10: Elbow plots. The elbow plots that result when clustering data from (a) Dataset 1 using the Cell counts DRDV, (b) Dataset 1 using the Final cell density DRDV, (c) Dataset 2 using the Cell counts DRDV, (d) Dataset 2 using the Final cell density DRDV
@@
Figure 11: Dataset 1 Analysis using the Final cell density DV. (a) The ($\eta_{1},c_{a}$) Parameter partition plot for Dataset 1 when using the Cell counts DV. Snapshots at times $t=2,6,\text{ and }10$ days from the representative model simulations from (b-d) Cluster 1, (e-g) Cluster 2, (h-j) Cluster 3, and (k-m) Cluster 4.
Figure 12: Confusion matrices. The confusion matrices that result when performing clustering robustness assessment for (a) Dataset 1 using the Cell counts DRDV, (b) Dataset 1 using the Final cell density DRDV, (c) Dataset 2 using the Cell counts DRDV, (d) Dataset 2 using the Final cell density DRDV
@@
Figure 13: Represenative DVs for Dataset1. Representative Cell counts DVs for Dataset 1 from (a-d) clusters 1-4. Representative Final cell density DVs for Dataset 1 from (e-h) clusters 1-4.
Figure 14: Represenative DVs for Dataset2. Representative Cell counts DVs for Dataset 2 from (a-d) clusters 1-4. Representative Final cell density DVs for Dataset 2 from (e-h) clusters 1-4.
@@
Figure 15: Dataset 2 Analysis using the Final cell density DV. (a-d) Ridgeline plots for Clusters 1-4. Each parameter $p$ is log-normalized as $\log_{2}(p/p_{base})$, where $p_{base}$ denotes the baseline model value from Table 1. Snapshots at times $t=2,6,\text{ and }10$ days from the representative model simulations from (e-g) Cluster 1, (h-j) Cluster 2, (k-m) Cluster 3, and (n-p) Cluster 4.
@@
$c_{a}$ $<10^{-3}$ $<10^{-3}$ $<10^{-3}$ $<10^{-3}$ $<10^{-3}$ $<10^{-3}$ Sensitive
$c_{d}$ $<10^{-3}$ $0.008$ $<10^{-3}$ $<10^{-3}$ $<10^{-3}$ $<10^{-3}$ Sensitive
$c_{m}$ 0.307 0.626 0.995 $<10^{-3}$ 0.098 0.712 Insensitive
@@
$\eta_{3}$ $<10^{-3}$ $<10^{-3}$ 1.0 $<10^{-3}$ $<10^{-3}$ $<10^{-3}$ Sensitive
Table 5: Determining sensitive Dataset 2 parameters from the Final cell density DRDV. We designate $p$-values below 0.05 as being significant. If a parameter receives $p$-values below 0.05 for the majority of cluster pairs, then we classify the parameter as sensitive.

### v3

@@ 4 Discussion and conclusions
Many methodologies exist to perform SA; two of the most widely-used approaches include the Morris and Sobol’ Methods [morris_factorial_1991, sobol_sensitivity_1990, sobol_sensitivity_1993, ten_broeke_which_2016]. We compare the benefits and drawbacks of SSRCA with these approaches in Table 4. All three methods can determine sensitive parameters, however, the Sobol’ Method fails to identify common model patterns because it computes sensitivity indices by aggregating results over the entire parameter space. The Morris method can identify common model patterns by varying one parameter at a time and tracking how the outputs change. However, the Morris method cannot perform extensive feature mapping for these patterns because it ignores parameter interactions. A strength of the Sobol’ Method is that it accounts for these interaction effects when computing its sensitivity indices. We find that the SSRCA Method can both identify model patterns and perform exhaustive feature mapping. Furthermore, we can vary multiple parameters at once to incorporate interaction effects into this analysis. A key benefit of the Morris method is its computational efficiency; previous studies estimate that the cost of the Morris method scales as $\sim 10$ times the number of parameters, whereas the cost of the Sobol’ Method scales as $\sim 10^{3}$ times the number of parameters [bergman_efficient_2025, morris_factorial_1991, marino_methodology_2008]. As a first pass, we implemented SSRCA on a dense set of parameters where we varied 10 parameters, 2 at a time. The SSRCA Method can be used with any parameter sampling, however, and in future work we plan to assess how the number of required simulations scales with the number of parameters. We will further examine its use with more efficient parameter sampling techniques, including Latin Hypercube sampling [marino_methodology_2008].
@@
Identify common model
patterns
yes no yes
@@
Computationally efficient yes no
further investigation
required
@@
Another strength of SSRCA is its flexibility with the choice of methodologies used within the pipeline. We found similar results in this work whether we used the Cell counts DV or Final Cell Density DV. This is in contrast to the Sobol’ Method, whose sensitivity analysis results were significantly different when using the Cell counts or Final cell density DVs, even though the underlying model simulations were unchanged. This robustness to DV choice is crucial, as ABMs are broadly used for biological problems, and the appropriate DV choices depends on the application at hand. For example, many studies rely on order parameters to characterize ABM behavior; alternatively, topological data analysis (TDA) is a recent field of research that has found success in summarizing many types of ABMs with applications in complex systems [bhaskar_analyzing_2019, bhaskar_topological_2023, gharooni-fard_computational_2024]. SSRCA can be used with any of these DVs, as it only requires they are uniform length. Additionally, we used principal components analysis and $k$-means clustering for dimensionality reduction and clustering, respectively, in this work, but any suitable algorithm can be used for these processes.
@@
We applied the SSRCA Methodology to a 2-dimensional ABM of tumor spheroid growth and identified the $c_{a}$, $c_{d}$, $\eta_{1}$, $\eta_{3}$ parameters as the most sensitive ones. These parameters make biological sense as sensitive parameters because they control cell death and entry into the cell cycle. These results may inform the future use of tumor spheroid models to interpret experimental data by focusing on these processes. For example, the 3-dimensional Klowss Model was calibrated to experimental data from human primary melanoma cells [4D_Tumor_Model]; modeling similar data from other cell lines will likely require different parameter values. Our work proposes that thoroughly exploring all unknown parameters is not necessary; instead, one may more efficiently vary the four identified sensitive parameters to fit the model to new data. Our proposed methodology may be of use for other modeling frameworks. For example, the authors of [jin_mathematical_2021] could not perform parameter estimation due to their differential equation model containing 19 unknown parameters. The SSRCA Methodology could be applied to this model to learn an informative subset of its parameters for parameter estimation.
@@
There are limitations associated with the SSRCA algorithm and our implementation of it in this work. While we can use the SSRCA Methodology to identify sensitive parameters, it is currently not able to rank parameters based on their sensitivity. We plan to further analyze the model and/or develop metrics that can be used in combination with SSRCA to enable such parameter ranking. We also demonstrated the utility of the SSRCA Method when applied to informative DVs, however, it has been reported previously that some DVs are more informative than others depending on the context [bhaskar_analyzing_2019, angio_tda, nardini_statistical_2023, topaz_topological_2015]. It would be interesting to explore if the SSRCA algorithm can be used to compare the information content between various DVS, or even to propose an optimal DV choice. This may be especially relevant when applying the SSRCA Method to more complex ABMs in the future. We applied the SSRCA Methodology to a simple radially symmetric and two-dimensional tumor spheroid ABM in this work; these simplifications yield simulations with similar output DVs when parameter values are kept constant. Applying our proposed methods to more complicated ABMs (e.g., radially asymmetric and/or 3-dimensional models) in the future will lead to more complex and noisy DVs, and it is not clear how SSRCA will perform in the face of these challenges or if the most sensitive parameters will be changed. We plan to perform consistency analyses to determine how many model simulations from each parameter combination are necessary to mitigate these increased uncertainly levels [gerstman_uncertainty_2020]. It will also be of interest to apply the SSRCA Methodology to ABMs from other biological contexts, including cell biology, epidemiology, and ecology.
Data and code availability: All code to simulate this study’s data and perform data analysis are publicly available at https://github.com/e-rohr/FUCCI_ABM2D.
@@
Appendix S1 Model overview
We simulate an agent-based model of tumor spheroid growth developed in [4D_Tumor_Model]. This model is hybrid, meaning we track both discrete agent locations and a continuous nutrient concentration (e.g., oxygen) within the spheroid. Tumor cells are represented as discrete agents centered at $\textbf{x}_{n}(t)=(x_{n}(t),y_{n}(t))$ for $n=1,\ldots,N(t)$, where $N(t)$ is the total number of agents at time $t$. The spatio-temporal distribution of the nutrient concentration is non-dimensionalized and represented by $c(\textbf{x},t)\in[0,1]$. In addition to agent locations, we also model how cells progress through the cell cycle. The agents are colored according to FUCCI: agents in the G1 phase are colored red, agents in the early S (eS) phase are colored yellow, and agents in the G2/S/M phase are colored green. Following [4D_Tumor_Model], we assume the non-dimensional, local nutrient concentration, $c(\textbf{x},t)$, regulates cell motility, proliferation, and death [greenspan_models_1972, Haass, ward_mathematical_1997]. The model is simulated in a cubic domain, $\Omega$, with side length $1000~\mu\mathrm{m}$, which we chose to mitigate computational costs while ensuring that agents do not reach the domain boundary during simulation.
We now describe how tumor cells in the model spheroid undergo mitosis, cell motility, and cell death, and then detail how the nutrient spreads and is consumed. All model parameters are described in Table 1. The tumor cell processes are modelled as discrete events using the Gillespie Algorithm [Gillespie]. We initialize the ABM in a similar manner as [4D_Tumor_Model], where initial cell colors are randomized to produce proportions of red, yellow, and green cells that match experimental observations.
S1.1 Agent Dynamics
@@
$m(c)$ $=(m_{max}-m_{min})\frac{c^{\eta_{2}}}{c_{m}^{\eta_{2}}+c^{\eta_{2}}}+m_{min},$ (S7)
$d(c)$ $=(d_{max}-d_{min})\left(1-\frac{c^{\eta_{3}}}{c_{d}^{\eta_{3}}+c^{\eta_{3}}}\right)+d_{min},$ (S8)
where all parameters are described in Table 1. When migrating, an agent is displaced $\mu$ units in the randomly chosen direction. Following [4D_Tumor_Model], cellular interactions are ignored during cell migration, cell cycle progression and death; i.e., these process are unaffected by cell mechanics, crowding, and/or cell-cell interactions. When an agent dies, it is removed from the simulation by setting its status to dead, and its final location is recorded.
S1.2 Nutrient dynamics
@@
$\frac{\partial C}{\partial t}=D\nabla^{2}C-\kappa Cv,\text{ in }\Omega,$ (S9)
@@
with diffusivity $D>0~[\mu\mathrm{m}^{2}~\mathrm{h}^{-1}]$, consumption rate $\kappa>0~[\mu m^{2}~(\mathrm{h}~\text{cells})^{-1}]$, and where $v(\textbf{x},t)\geq 0~[\text{cells }\mu m^{-2}]$ is the density of agents at position x and time $t.$ On the boundary, $\partial\Omega$, a maximum far-field concentration $C(\textbf{x},t)=c_{b}$ is imposed. We non-dimensionalize this independent variable by setting $c(\textbf{x},t)=C(\textbf{x},t)/c_{b}$, so that $c=1$ on $\partial\Omega$, and $c(\textbf{x},t)\in[0,1]$. We approximate the reaction-diffusion Equation (S9) using the quasi-steady state equation
$0=\nabla^{2}c-\alpha cv,\text{ in }\Omega,$ (S10)
@@
where $\alpha=\kappa/D>0\ [\mu\mathrm{m}~\mathrm{cell}^{-1}]$. During simulations, the spatio-temporal distribution of the chemical is modelled by solving Equation (S10) $t^{*}$ units using a finite volume method on a uniform structured mesh with node spacing $h$. Following [4D_Tumor_Model], we use $t^{*}=1$. The cell density $v(\textbf{x},t)$ is approximated by setting $v(\textbf{x}_{i,j},t)=N_{i,j}/h^{2}$, where $N_{i,j}$ is the number of agents within the control volume surrounding the node located at $(x_{i},y_{j})$ and $h^{2}$ is the volume of the control volume.
S1.3 Initializing the ABM
[…]
Appendix S2 Interpretation of ridgeline plots
In this section, we discuss how to interpret sensitive and insensitive parameters from ridgeline plots over various data clusters. We created a simple dataset by performing 500 samples for two parameters, $x_{1}\text{ and }x_{2}$, from a uniform distribution on the domain $[0,1]$. We will illustrate how to identify (a) sensitive vs. insensitive parameters and (b) parameter interactions through two example clusterings of this data. In clustering 1, We designate each sample $(x_{1},x_{2})$ according to the following rule: samples with $x_{2}\in[0,1/3]$ belong to cluster 1, samples with $x_{2}\in(1/3,2/3]$ belong to cluster 2, and samples with $x_{2}\in(2/3,1]$ belong to cluster 3 (Supplementary Figure S8(a)). Here, only $x_{2}$ determines a sample’s clustering assignment (so it is sensitive) and $x_{1}$ has no impact (so it is insensitive). From this data and clustering, the ridgeline plot is constructed as follows. For each cluster, we determine all values for $x_{1}$ and $x_{2}$ within that cluster and plot these marginal distributions for each parameter (Supplementary Figure S8(b-d)). The sensitivity of $x_{2}$ is shown through the three peaked distributions for $x_{2}$ in the ridgeline plots. The insensitivity of $x_{1}$ is shown through the three widely-distributed distributions for $x_{1}$ in the ridgeline plots.
We created data clustering 2 to demonstrate how ridgeline plots show parameter interactions between clusters. In this clustering, we place samples where $x_{2}-x_{1}\leq-1/3$ into cluster 1, samples where $1/3\leq x_{2}-x_{1}\leq 1/3$ into cluster 2, and samples where $x_{2}-x_{1}\leq 1/3$ into cluster 3 (Supplementary Figure S8(e)). Here, the parameters interact in generating the clusterings, as cluster 1 contains samples with large values of $x_{1}$ and small values of $x_{2}$, cluster 2 contains samples with all values of $x_{1}$ and $x_{2}$, and cluster 3 contains samples with small values of $x_{1}$ and large values of $x_{2}$. These various distributions are represented in the ridgeline plots for all clusters (Supplementary Figure S8(f-h)).
Figure S6: Determining the baseline $N(0)$ value. To determine the baseline value of $N(0)$, we visually compared simulations of the 2-dimensional Klowss model (colored lines) for various values of $N(0)$ to the $z=0$ cross section of the 3-dimensional Klowss Model from [4D_Tumor_Model] (black squares) for cells in the (a) G1 state, (b) Early S state, (c) S/G2/M state, and (d) dead state .
Figure S7: Determining the baseline $\alpha$ value. To determine the baseline value of $\alpha$, we visually compared simulations of the 2-dimensional Klowss model (colored lines) for various values of $\alpha$ to the $z=0$ cross section of the 3-dimensional Klowss Model from [4D_Tumor_Model] (black squares) for cells in the (a) G1 state, (b) Early S state, (c) S/G2/M state, and (d) dead state .
Figure S8: Ridgeline interpretation plot. (a) Uniformly sampled data $(x_{1},x_{2})$ where only $x_{2}$ is sensitive to the cluster labeling. (b)-(d) Ridgeline plots for Clusters 1-3 for this labeling. (e) Uniformly sampled data $(x_{1},x_{2})$ where $x_{1}$ and $x_{2}$ interact for the data cluster labeling. (f)-(h) Ridgeline plots for Clusters 1-3 for this labeling.
@@
Figure S9: Scree plots. The scree plots that result when performing dimensionality reduction using (a) the small 2-parameter dataset using the Cell counts DRDV, (b) the small 2-parameter dataset using the Final cell density DRDV, (c) the large 10-parameter dataset using the Cell counts DRDV, (d) the large 10-parameter dataset using the Final cell density DRDV
@@
Figure S10: Elbow plots. The elbow plots that result when clustering data from (a) the small 2-parameter dataset using the Cell counts DRDV, (b) the small 2-parameter dataset using the Final cell density DRDV, (c) the large 10-parameter dataset using the Cell counts DRDV, (d) the large 10-parameter dataset using the Final cell density DRDV
@@
Figure S11: The small 2-parameter dataset Analysis using the Final cell density DV. (a) The ($\eta_{1},c_{a}$) Parameter partition plot for the small 2-parameter dataset when using the Cell counts DV. Snapshots at times $t=2,6,\text{ and }10$ days from the representative model simulations from (b-d) Cluster 1, (e-g) Cluster 2, (h-j) Cluster 3, and (k-m) Cluster 4.
@@
Figure S12: Represenative DVs for the small 2-parameter dataset. Representative Cell counts DVs for the small 2-parameter dataset from (a-d) clusters 1-4. Representative Final cell density DVs for the small 2-parameter dataset from (e-h) clusters 1-4.
@@
Figure S13: The large 10-parameter dataset Analysis using the Final cell density DV. (a-d) Ridgeline plots for Clusters 1-4. Each parameter $p$ is log-normalized as $\log_{2}(p/p_{base})$, where $p_{base}$ denotes the baseline model value from Table 1. Snapshots at times $t=2,6,\text{ and }10$ days from the representative model simulations from (e-g) Cluster 1, (h-j) Cluster 2, (k-m) Cluster 3, and (n-p) Cluster 4.
Figure S14: Represenative DVs for the large 10-parameter dataset. Representative Cell counts DVs for the large 10-parameter dataset from (a-d) clusters 1-4. Representative Final cell density DVs for the large 10-parameter dataset from (e-h) clusters 1-4.
@@
$c_{a}$ $10^{-3} <$$10^{-3} <$$10^{-3} <$$10^{-3} <$$10^{-3} <$$10^{-3}$ Sensitive
$c_{d}$ $10^{-3}$ $0.008$ $10^{-3} <$$10^{-3} <$$10^{-3} <$$10^{-3}$ Sensitive
$c_{m}$ 0.307 0.626 0.995 $10^{-3}$ 0.098 0.712 Insensitive
@@
[…]
Table S5: Determining sensitive the large 10-parameter dataset parameters from the Final cell density DRDV. We designate $p$-values below 0.05 as being significant. If a parameter receives $p$-values below 0.05 for the majority of cluster pairs, then we classify the parameter as sensitive.